# LDMol: Text-to-Molecule Diffusion Model with Structurally Informative Latent Space

### v2

@@ Abstract
With the emergence of diffusion models as the frontline of generative models, many researchers have proposed molecule generation techniques with conditional diffusion models. However, the unavoidable discreteness of a molecule makes it difficult for a diffusion model to connect raw data with highly complex conditions like natural language. To address this, we present a novel latent diffusion model dubbed LDMol for text-conditioned molecule generation. LDMol comprises a molecule autoencoder that produces a learnable and structurally informative feature space, and a natural language-conditioned latent diffusion model. In particular, recognizing that multiple SMILES notations can represent the same molecule, we employ a contrastive learning strategy to extract feature space that is aware of the unique characteristics of the molecule structure. LDMol outperforms the existing baselines on the text-to-molecule generation benchmark, suggesting a potential for diffusion models can outperform autoregressive models in text data generation with a better choice of the latent domain. Furthermore, we show that LDMol can be applied to downstream tasks such as molecule-to-text retrieval and text-guided molecule editing, demonstrating its versatility as a diffusion model.
@@ 1 Introduction
Figure 1: Different strategies of data domain selection for molecule diffusion models. (a) The model directly learns the raw representation of the molecule such as string format, point cloud coordinates, etc. (b) An autoencoder can be employed to let the generative model learn its latent distribution. (c) A regularized, chemically pre-trained encoder can provide a latent space learnable by external generative models.
@@
However, a discrepancy between molecule data and common data domains like images makes it hard to connect the diffusion models to molecule generation. Whereas diffusion models are deeply studied on a continuous data domain with gaussian noise, any molecule modality has inevitable discreteness such as atom and bond type, connectivity, and SMILES tokens as shown in Figure 1-(a). Naively converting discrete values into real numbers would ignore rich correlation within the components, and still retain certain restrictions on the data. As a result, diffusion models trained on raw molecule data often failed to faithfully follow the given conditions or showed poor data quality (e.g., low validity) as the condition became more sophisticated like natural language. Most molecule diffusion models presented so far have used a few, relatively simple conditions to control, while major developments in text-conditioned molecule generative models were based on autoregressive models.
@@
To overcome the limitation of the previous works (Xu et al., 2023) that mainly focused on resolving the discreteness of the molecular domain space by training with reconstruction loss (Figure 1-(b)), we aim to build a latent space encoder to extract rich and refined information about the molecule structure such as Figure 1-(c). Specifically, we design a novel Latent Diffusion Molecular generative model (LDMol) for text-conditioned molecule generation, trained on the latent space of the separately pre-trained molecule encoder. By preparing an encoder to provide a chemically useful and interpretable feature space, our model can more easily connect the molecule data with the highly complicated condition of natural texts. In the process, we suggest a novel contrastive learning strategy of training a molecular encoder to encode a structural invariant of the molecule.
@@
The experimental results show that LDMol can outperform many state-of-the-art autoregressive models and generate valid SMILES that meet the input text condition. Considering SMILES as a variation of text data, we report one of the first diffusion models that successfully surpassed autoregressive models in text data generation. This suggests the possibility of improving existing diffusion models (Lovelace et al., 2024) for natural language through careful design of the latent space. Furthermore, LDMol can leverage the learned score function and be applied to several multi-modal downstream tasks such as molecule-to-text retrieval and text-guided molecule editing, without additional task-specific training. We summarize the contribution of this work as follows:
@@
  • LDMol outperforms the text-to-molecule generation baselines, and its modeled conditional score function allows LDMol to retain the advanced attributes of diffusion models including various applications like molecule-to-text retrieval and text-guided molecule editing.
@@ 2 Background
$\theta^{*}$ $=\text{arg}\min_{\theta}\mathbb{E}_{x_{0},y,t,\epsilon}||\epsilon-\epsilon_{\theta}(x_{t},t,c)||_{2}^{2}$ (5)
Diffusion models can perform their diffusion process on the latent data space instead of the raw data when the raw data is too complex or high-dimensional (Vahdat et al., 2021; Rombach et al., 2022). Specifically, using a pre-trained pair of an encoder $\mathcal{E}(\cdot)$ and its corresponding decoder $\mathcal{D}(\cdot)$, datapoint $x$ is encoded into $z:=\mathcal{E}(x)$ that is used to train the diffusion model. Generated latent $z^{\prime}$ from the pre-trained diffusion model is transformed back to $x^{\prime}:=\mathcal{D}(z^{\prime})$ on the original data domain. The latent diffusion model was also adopted for discrete texts (Lovelace et al., 2024), where the latent space was constructed by a denoising autoencoder.
Conditional molecule generation. As a promising tool for many important chemical and engineering tasks like de novo drug discovery and material design, conditional molecule generation has been extensively studied with various models including recurrent neural networks (Segler et al., 2017), graph neural networks (Lim et al., 2020), and variational autoencoders (Gómez-Bombarelli et al., 2018; Lim et al., 2018). With the advent of large and scalable pre-trained models with transformers (Vaswani et al., 2017), the controllable conditions became more abundant and complicated (Bagal et al., 2021; Chang & Ye, 2024), and reached a text-guided molecule generation (Edwards et al., 2022; Su et al., 2022; Liu et al., 2024a) which requires a deep comprehension ability for natural language.
Recent works attempted to import the success of the diffusion model into molecule generation. Several graph-based and point cloud-based works have built conditional diffusion models that could generate molecules with simple chemical and biological conditions (Hoogeboom et al., 2022; Luo et al., 2023; Trippe et al., 2023). Gong et al. (2024) attempted a text-conditioned molecule diffusion model trained on the sequence of tokenized SMILES indices. However, these models treated discrete molecule attributes with continuous Gaussian diffusion models, introducing arbitrary numeric values and suboptimal performances. Xu et al. (2023) employed an autoencoder to build a diffusion model on a smooth latent space, but its controllable conditions were still limited to several physiochemical properties.
@@ 3 Methods
In this section, we explain the overall model architecture and training procedure of the proposed LDMol, which are briefly illustrated in Figure 2. We first discuss the conditions the molecule encoder should meet and present our strategy to construct an encoder that can extract structural information from the raw molecule representation. Then we describe how we prepared a complete encoder-decoder pair with this encoder and trained a text-conditioned diffusion model on its latent space.
@@
3.1 Chemically informative SMILES autoencoder
@@
The primary goal of introducing autoencoders for image latent diffusion models is to map raw images into a low-dimensional space which reduces the computation cost (Vahdat et al., 2021; Rombach et al., 2022). This is plausible because a high-resolution image has an enormous dimension on a pixel domain, yet each pixel contains little information.
@@
In this work for molecule generation, we utilized SMILES, one of the most widely used molecule representations in chemical databases and benchmarks, and built a SMILES encoder to map raw SMILES strings into a latent vector. In this case, the role of our SMILES autoencoder has to be different from that of the autoencoders for images; a molecule structure can be fully expressed by only a sequence of $L$ integers for SMILES tokens, where $L$ is the maximum token length. However, each token carries significant information, and hidden interactions between these tokens are much more complicated than interactions between image pixels. Therefore, the SMILES encoder should focus more on extracting chemical meaning into the latent space, even if it results in a latent space with more dimension than the raw SMILES string.
@@
A number of molecule encoders (Wang et al., 2019; Liu et al., 2024a; Zeng et al., 2022; Liu et al., 2023a) have been presented that can extract various useful chemical features including biochemical activity or human-annotated descriptions. Nonetheless, these molecule encoders aim to extract certain desired features rather than encode all the information about the molecule structure. Therefore the input cannot be fully restored from the model output.
@@
Although autoencoders with appropriate regularization (e.g., KL-divergence loss (Kingma, 2013; Gómez-Bombarelli et al., 2018)) provide a continuous and reconstructible molecular latent space, their encoder output is not guaranteed to possess the characteristic of the underlying molecular structure, beyond the minimal information to reconstruct the input string. To visualize this, we prepared a trained $\beta$-VAE (Higgins et al., 2017) and measured the feature distance between two SMILES from the same molecule obtained via SMILES enumeration (Bjerrum, 2017). Here, SMILES enumeration is the process of writing out all possible SMILES of the same molecule, as illustrated in Figure  3-(a). Figure  3-(b) shows that $\beta$-VAE had difficulties assimilating features from the same molecule compared to the one between random SMILES pairs, indicating that it couldn’t capture the intrinsic features beneath the SMILES string. This inherited inconsistency makes it difficult for later models that learn this latent space to figure out the connection between the latent and the molecule, which could eventually degrade the performance of the generative model with complex conditions. If a molecule encoder can truly provide a unique representation of the atoms and their connectivity, all SMILES representing the same molecule must be encoded into the same output and should be differentiated from the other random SMILES.
@@
Accordingly, here we propose three conditions that our SMILES autoencoder’s latent space has to satisfy: enable reconstruction of the input, have as small dimensions as possible, and embed chemical information that can be readily learned by diffusion models.
@@
Encoder design. In this respect, we trained our SMILES encoder with contrastive learning (Figure 2-(a)), which aims to learn better representation by assimilating features containing similar information (i.e. positive pair) and distancing semantically unrelated features (i.e. negative pair). We defined two enumerated SMILES from the same molecule as a positive pair and two SMILES from different molecules as a negative pair.
@@
Here, we argue that the proposed contrastive learning with SMILES enumeration can train the encoder to encapsulate the structural information of the input molecule. Contrastive learning learns invariant for the augmentations applied on positive pairs (Zhang & Ma, 2022), and it is known that a good augmentation should reduce as much mutual information between positive pairs as possible while preserving relevant information (Tian et al., 2020). Meanwhile, enumerated SMILES of the same molecule are obtained by traversing the nodes and edges in the molecular graph with a different visiting order. Therefore, to detect all possible enumerated SMILES and find SMILES-enumeration-invariant, the model has to understand the entire connectivity between atoms. This makes the encoder output a unique characteristic that captures the overall molecule structure. Compared to the hand-crafted augmentations previously presented for molecule contrastive learning (You et al., 2020), enumerated SMILES pairs have less mutual information since we utilize all possible variations in the SMILES format. And since all enumerated SMILES are guaranteed to represent an identical molecule, there is no relevant information loss during the augmentation. Figure  3-(a) confirms that our LDMol trained with the contrastive learning with SMILES enumeration now correctly assimilates features from the same molecule compared to the one between random SMILES pairs.
@@
Specifically, a SMILES string is fed into the SMILES encoder $\mathcal{E}(\cdot)$ with a special “[SOS]” token which denotes the start of the sequence. For a batch of $N$ input SMILES $M=\{m_{1},m_{2},\dots,m_{k},\dots,m_{N}\}$, we prepare a positive pair SMILES $m_{k}^{\prime}$ for each $m_{k}$ by SMILES enumeration to construct $M^{\prime}=\{m_{1}^{\prime},m_{2}^{\prime},\dots,m_{k}^{\prime},\dots,m_{N}^{\prime}\}$. After $M$ and $M^{\prime}$ are passed through the SMILES encoder, we fed each SMILES’ output vector corresponding to the [SOS] token into an additional linear projection and normalization layer, denoting its output as $v_{k}$ and $v_{k}^{\prime}(k=1,2,\dots,N)$. Assimilating $v_{k}$ and $v_{k}^{\prime}$ from the positive pairs and distancing the others can be done by minimizing the following InfoNCE loss (Chen et al., 2020).
$\mathcal{L}_{contrastive}(M,M^{\prime})=-\sum_{k=1}^{N}\log\frac{\exp(v_{k}\cdot v_{k}^{\prime}/\tau)}{\sum_{i=1}^{N}\exp(v_{k}\cdot v_{i}^{\prime}/\tau)}\;$ (6)
@@
$\mathcal{L}_{enc}(M,M^{\prime})=\mathcal{L}_{contrastive}(M,M^{\prime})+\mathcal{L}_{contrastive}(M^{\prime},M)\;$ (7)
While our encoder was pre-trained to map structurally similar molecules to similar vectors, we found that the encoder tried to map stereoisomers to almost the same point in the latent space; indeed, in terms of structural differences, stereoisomers are the most challenging molecules to distinguish. To resolve this, if the training data has a stereoisomer, we consider it as “hard-negative” samples and include it in the loss calculation batch with the original data.
Compressing the latent space. The pre-trained SMILES encoder maps a molecule into a vector of size $[L\times d_{enc}]$, where $d_{enc}$ is a feature size of the encoder. To avoid the curse of dimensionality and construct a more learnable feature space for diffusion models, we additionally employed a linear compression layer $f(\cdot)$ (Figure 2-(b)) to reduce the dimension from $[L\times d_{enc}]$ to $[L\times d_{z}]$. The range of this linear layer output is a target domain of our latent diffusion model.
Decoder design. When a SMILES $m$ is passed through the SMILES decoder and the compression layer, the SMILES decoder reconstructs $m$ from $f(\mathcal{E}(m))$. Following many major works that treated SMILES as a variant of language data (Segler et al., 2017; Chang & Ye, 2024; Ross et al., 2022), we built an autoregressive transformer (Vaswani et al., 2017) as our decoder (Figure 2-(b)) which is widely used to successfully generate sequential data (Brown et al., 2020). Specifically, starting from the [SOS] token, the decoder predicts the next SMILES token using information from $f(\mathcal{E}(m))$ with cross-attention layers. When $\{t_{0},t_{1},\dots,t_{n}\}$ is the token sequence of $m$, the decoder is trained to minimize the next-token prediction loss described as Eq. (8). Here, the decoder and the compression layer are jointly trained while the encoder’s parameter is frozen.
@@
As long as the encoder output contains complete information of the molecule structure, the decoder can reconstruct the frozen encoder input. After being fully trained, the decoder was able to reconstruct roughly 98% of the SMILES encoder input.

### v4

# LDMol: A Text-to-Molecule Diffusion Model with Structurally Informative Latent Space Surpasses AR Models

@@ Abstract
With the emergence of diffusion models as a frontline generative model, many researchers have proposed molecule generation techniques with conditional diffusion models. However, the unavoidable discreteness of a molecule makes it difficult for a diffusion model to connect raw data with highly complex conditions like natural language. To address this, here we present a novel latent diffusion model dubbed LDMol for text-conditioned molecule generation. By recognizing that the suitable latent space design is the key to the diffusion model performance, we employ a contrastive learning strategy to extract novel feature space from text data that embeds the unique characteristics of the molecule structure. Experiments show that LDMol outperforms the existing autoregressive baselines on the text-to-molecule generation benchmark, being one of the first diffusion models that outperforms autoregressive models in textual data generation with a better choice of the latent domain. Furthermore, we show that LDMol can be applied to downstream tasks such as molecule-to-text retrieval and text-guided molecule editing, demonstrating its versatility as a diffusion model.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Different strategies of data domain selection for molecule diffusion models. (a) The model directly learns the raw representation of the molecule such as string tokens. (b) An autoencoder can be employed to let the generative model learn its latent distribution. (c) A regularized, chemically pre-trained encoder can provide a latent space readily learnable by external generative models.
@@
However, a discrepancy between molecule data and common data domains like images makes it hard to connect the diffusion models to molecule generation. Whereas diffusion models are deeply studied on a continuous data domain with Gaussian noise, any molecule modality has inevitable discreteness such as atom and bond type, connectivity, and SMILES tokens (Figure 1-(a)). As a result, diffusion models trained on raw molecule data often failed to faithfully follow the given conditions or showed poor data quality (e.g., invalid molecules) as the condition became more sophisticated like natural language. Most molecule diffusion models presented so far have used a few, relatively simple conditions to control, while major developments in text-to-molecule generative models were based on autoregressive models.
@@
To overcome this gap, we suggest that a latent domain (Vahdat et al., 2021; Rombach et al., 2022) is essential to train effective diffusion models for complex molecule generation tasks. Moreover, beyond the limitation of the previous works (Xu et al., 2023) that mainly focused on resolving the discreteness with naive reconstruction loss (Figure 1-(b)), we report that a latent encoder extracting rich and refined information about the molecule structure can further improve the generative model performance (Figure 1-(c)). Specifically, we design a novel Latent Diffusion Molecular generative model (LDMol) for text-conditioned molecule generation, trained on the latent space of the separately pre-trained molecule encoder. By preparing an encoder to provide a chemically useful and interpretable feature space, our model can more easily connect the molecule data with the highly complicated condition of natural text. In the process, we suggest a novel contrastive encoder training strategy by minimizing mutual information between positive SMILES pairs to encode a unique structural characteristic.
@@
Extensive experimental results show that LDMol can outperform many state-of-the-art autoregressive models and generate valid SMILES that meet the input text condition. Considering SMILES as a variation of text data, we report one of the first diffusion models that successfully surpassed autoregressive models in textual data generation. This may suggest the possibility of improving existing diffusion models (Lovelace et al., 2024) for natural language through careful design of the latent space. Furthermore, LDMol can leverage the learned score function and be applied to several multi-modal downstream tasks such as molecule-to-text retrieval and text-guided molecule editing, without additional task-specific training. We summarize the contribution of this work as follows:
@@
  • LDMol outperforms the text-to-molecule generation baselines, and its modeled conditional score function enables the advanced attributes of diffusion models including various applications like molecule-to-text retrieval and text-guided molecule editing.
@@ 2 Background
$\theta^{*}$ $=\text{arg}\min_{\theta}\mathbb{E}_{x_{0},c,t,\epsilon}||\epsilon-\epsilon_{\theta}(x_{t},t,c)||_{2}^{2}$ (5)
After the success of diffusion models in the image and video domain, various works tried to build diffusion models to generate text data. While many suggested training diffusion models on text tokens (Austin et al., 2021), word embedding (Li et al., 2022), or text autoencoder latent space (Lovelace et al., 2024), their performance has been suboptimal compared to autoregressive models (Brown et al., 2020). We assume that this can be resolved with a better latent space design that reflects the characteristics of the data domain, and suggest a latent diffusion model that outperforms autoregressive models for textual data.
Conditional molecule generation. As a promising tool for many important chemical and engineering tasks like de novo drug discovery and material design, conditional molecule generation has been extensively studied with various models including recurrent neural network (RNN)s (Segler et al., 2017), bidirectional RNN (Grisoni et al., 2020), graph neural networks (Lim et al., 2020), and variational autoencoders (Gómez-Bombarelli et al., 2018; Lim et al., 2018). With the advent of large and scalable pre-trained models with transformers (Vaswani et al., 2017), the controllable conditions became more abundant and complicated (Bagal et al., 2021; Chang & Ye, 2024). Recent works reached a text-guided molecule generation (Edwards et al., 2022; Su et al., 2022; Liu et al., 2024a) leveraging a deep comprehension ability for natural language, especially with recent emergence of large language model (LLM)s (Liu et al., 2024b).
Recent works attempted to import the success of the diffusion model into molecule generation. Several graph-based and point cloud-based works have built conditional diffusion models that could generate molecules with simple chemical and biological conditions (Hoogeboom et al., 2022; Luo et al., 2023; Trippe et al., 2023). Gong et al. (2024) attempted a text-conditioned molecule diffusion model trained on the sequence of tokenized SMILES indices. However, these models treated discrete molecules with continuous Gaussian diffusion, introducing arbitrary numeric values and suboptimal performances. Xu et al. (2023) employed an autoencoder to build a diffusion model on a smooth latent space, but its controllable conditions were still limited to several physiochemical properties.
@@ 3 Methods
In this section, we explain the overall model architecture and training procedure of the proposed LDMol, which are briefly illustrated in Figure 2.
@@
3.1 Extracting structure-aware SMILES latent space
@@
The primary goal of introducing autoencoders for image latent diffusion models is to map raw images into a low-dimensional space, which reduces the computation cost (Vahdat et al., 2021; Rombach et al., 2022). This is plausible because a high-resolution image has an enormous dimension in the pixel domain, yet each pixel contains little information.
@@
In this work for molecule generation, we utilize a string-based notation SMILES, one of the most popular molecule representations in text-molecule pair databases and benchmarks. We built a SMILES encoder to map raw SMILES strings into a latent vector. In this case, the role of our SMILES autoencoder has to be different from that of the autoencoders for images; a molecule structure can be fully expressed by only a sequence of $L$ integers for SMILES tokens, where $L$ is the maximum token length. However, each token carries significant information, and hidden interactions between these tokens are much more complicated than interactions between image pixels. Therefore, the SMILES encoder should focus more on extracting chemical meaning into the latent space, even if it results in a latent space with more dimensions than the raw SMILES string.
@@
A number of molecule encoders (Wang et al., 2019; Liu et al., 2024a; Zeng et al., 2022; Liu et al., 2023a) have been presented that can extract various useful chemical features, including biochemical activity or human-annotated descriptions. Nonetheless, these molecule encoders aim to extract certain desired features rather than encode all the information about the molecule structure. Therefore, the input cannot be fully restored from the model output.
@@
Although autoencoders with appropriate regularization (e.g., KL-divergence loss (Kingma, 2014; Gómez-Bombarelli et al., 2018)) provide a continuous and reconstructible molecular latent space, their encoder output is not guaranteed to possess the characteristic of the underlying molecular structure, beyond the minimal information to reconstruct the input string. To visualize this, we prepared a trained $\beta$-VAE (Higgins et al., 2017) and measured the feature distance between two SMILES from the same molecule obtained via SMILES enumeration (Bjerrum, 2017). Here, SMILES enumeration is the process of writing out all possible SMILES of the same molecule, as illustrated in Figure  3-(a). Figure  3-(b) shows that $\beta$-VAE had difficulties assimilating features from the same molecule compared to the one between random SMILES pairs, indicating that it couldn’t capture the intrinsic features beneath the SMILES string. This inconsistency makes it difficult for later models that learn this latent space to figure out the connection between the latent and the molecule, which could eventually degrade the performance as the condition gets more complex like natural texts. Assuming most of the controllable conditions has unavoidable correlation with the molecule structure, we insist that latent domain where the feature proximity is more structurally meaningful would benefit the conditional generative model.
@@
Accordingly, here we propose three conditions that our SMILES autoencoder’s latent space has to satisfy: enable reconstruction of the input, have as small dimensions as possible, and embed molecular structural information that can be readily learned by diffusion models.
@@
Encoder design. In this respect, we train our SMILES encoder with contrastive learning (Figure 2-(a)), which aims to learn better representation by assimilating features containing similar information (i.e. positive pair) and distancing semantically unrelated features (i.e. negative pair). We define two enumerated SMILES from the same molecule as a positive pair and two SMILES from different molecules as a negative pair.
@@
Here, we argue that the proposed contrastive learning with SMILES enumeration can train the encoder to encapsulate the unique structural characteristics of the input molecule: Contrastive learning learns an invariant for the augmentations applied on positive pairs (Zhang & Ma, 2022), and it is known that a good augmentation should reduce as much mutual information between positive pairs as possible while preserving relevant information (Tian et al., 2020). Meanwhile, enumerated SMILES of the same molecule are obtained by traversing the nodes and edges in the molecular graph with a different visiting order. Therefore, to detect all possible enumerated SMILES and find SMILES-enumeration-invariant, the model has to understand the entire connectivity between atoms. This makes the encoder output a unique characteristic that captures the overall molecular structure. Compared to the hand-crafted augmentations previously presented for molecule contrastive learning (You et al., 2020), enumerated SMILES pairs have minimal mutual information since we utilize all possible variations in the SMILES format. And since all enumerated SMILES are guaranteed to represent an identical molecule, there is no relevant information loss during the augmentation. Figure 3-(a) and Appendix B.1 demonstrate that our LDMol trained with contrastive learning with SMILES enumeration now correctly assimilates features from the same molecule, and its latent space captures meaningful structural information of the molecule.
@@
Specifically, a SMILES string is fed into the SMILES encoder $\mathcal{E}(\cdot)$ with a special “[SOS]” token which denotes the start of the sequence. For a batch of $N$ input SMILES $M=\{m_{1},m_{2},\dots,m_{k},\dots,m_{N}\}$, we prepare a positive pair SMILES $m_{k}^{\prime}$ for each $m_{k}$ by SMILES enumeration to construct $M^{\prime}=\{m_{1}^{\prime},m_{2}^{\prime},\dots,m_{k}^{\prime},\dots,m_{N}^{\prime}\}$. After $M$ and $M^{\prime}$ are passed through the SMILES encoder, we feed each SMILES’ output vector corresponding to the [SOS] token into an additional linear projection and normalization layer, denoting its output as $v_{k}$ and $v_{k}^{\prime}(k=1,2,\dots,N)$. Assimilating $v_{k}$ and $v_{k}^{\prime}$ from the positive pairs and distancing the others can be done by minimizing the following InfoNCE loss (Oord et al., 2018).
$\mathcal{L}_{con}(M,M^{\prime})=-\sum_{k=1}^{N}\log\frac{\exp(v_{k}\cdot v_{k}^{\prime}/\tau)}{\sum_{i=1}^{N}\exp(v_{k}\cdot v_{i}^{\prime}/\tau)}\;$ (6)
@@
$\mathcal{L}_{enc}(M,M^{\prime})=\mathcal{L}_{con}(M,M^{\prime})+\mathcal{L}_{con}(M^{\prime},M)\;$ (7)
Compressing the latent space. The pre-trained SMILES encoder maps a molecule into a vector of size $[L\times d_{enc}]$, where $d_{enc}$ is the feature size of the encoder. To avoid the curse of dimensionality and construct a more learnable feature space for diffusion models, we additionally employed a linear compression layer $f(\cdot)$ (Figure 2-(b)) to reduce the dimension from $[L\times d_{enc}]$ to $[L\times d_{z}]$. Here, we build our compression layer as simply as possible to prevent its output from deviating from the previous structure-aware and regulated features (See A.2 for further justification). The range of this linear layer output is a target domain of our latent diffusion model.
Decoder design. When a SMILES $m$ is passed through the SMILES decoder and the compression layer, the SMILES decoder reconstructs $m$ from $f(\mathcal{E}(m))$. Note that the decoder knows nothing about SMILES distribution or its correlation with natural texts, and any design would be acceptable as long as it recovers SMILES from its latent. Following many major works that treated SMILES as a variant of language data (Segler et al., 2017; Chang & Ye, 2024), we built an autoregressive transformer (Vaswani et al., 2017) as our decoder (Figure 2-(b)) which is widely used to successfully generate sequential data with varied length (Brown et al., 2020). Specifically, starting from the [SOS] token, the decoder predicts the next SMILES token using information from $f(\mathcal{E}(m))$ with cross-attention layers. When $\{t_{0},t_{1},\dots,t_{n}\}$ is the token sequence of $m$, the decoder is trained to minimize the next-token prediction loss described as Eq. (8). Here, the decoder and the compression layer are jointly trained while the encoder’s parameter is frozen. After being fully trained, the decoder was able to reconstruct roughly 98% of the SMILES encoder input.